The K_a of a monoprotic weak acid is 7.53 times 10
Answer #1

HA ---------------> H+ + A-

0.107 0 0

0.107-x x x

Ka = x^2 / 0.107 - x

solve the equation to get x

x = 0.0248

% of ionisation = x *100 / 0.107 = 23.1775 %